OpenStudy (anonymous):

The area of a parallelogram is A = bh. Solve for h (height) in this formula.

OpenStudy (callisto):

Weird question... h = A/b ...

OpenStudy (anonymous):

A.h = Ab , B.h = A/b , C. h = b/A , D.h = A - b ??????????????

OpenStudy (callisto):

I've given you the answer, why can't you match it?

OpenStudy (anonymous):

oh i didn't notice, B thanks i'm very tired i'm up asking questions to get better at this i have a test in the a.m

OpenStudy (callisto):

It's B... pay attention next time :|